# disposition

### v2

@@ 5.4 Scalarization results for the concept of proper efficiency in the sense of Henig
Lemma 5.36.
Assume that $Y$ is a real locally convex space, $K\subseteq Y$ is a nontrivial cone, and $\bar{D}:=Y\setminus(-\operatorname{int}\,D)$ for some $D\in\mathcal{D}(K)$. Assume that $B_{K}$ and $B_{\bar{D}}$ are normlike-bases for $K$ and $\bar{D}$, respectively, constructed with respect to a continuous seminorm $\psi$. Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{\bar{D}}^{0}$ is compact, and
$({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$
@@
Proof 5.37.
First, note that $\emptyset\neq K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$ for all $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ taking into account the continuity of $\psi$. By Proposition 4.15, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
$x^{*}(d)+\alpha\psi(d)\geq 0>x^{*}(-k)+\alpha\psi(k)\quad\mbox{for all }d\in\bar{D}\mbox{ and all }k\in K\setminus\{0\}.$ (5.3)
Any $x\in C^{>}_{\psi}(x^{*},\alpha)$ satisfies $0>x^{*}(-x)+\alpha\psi(x)$, hence $-x\notin\bar{D}$ in view of (5.3), and so $x\in\operatorname{int}\,D$.
Theorem 5.38.
Assume that $Y$ is a real locally convex space, and $K\subseteq Y$ is a nontrivial cone. Let $\bar{x}\in{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$, i.e., there is $D\in\mathcal{D}(K)$ such that $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,D)$. Suppose that $D$ is closed. Define $\bar{D}:=Y\setminus(-\operatorname{int}\,D)$. Assume that $B_{K}$ and $B_{\bar{D}}$ are normlike-bases for $K$ and $\bar{D}$, respectively, constructed with respect to a continuous seminorm $\psi$. Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{\bar{D}}^{0}$ is compact, and
$({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$
Then, the conclusion of Theorem 5.25 holds.
Proof 5.39.
By Lemma 5.36 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
@@
$\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,D)\subseteq{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha)).$
Finally, noting that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ by Lemma 2.11 ($7^{\circ}$), Proposition 5.12 yields the rest of the conclusion.
Theorem 5.40.
Assume that $Y$ is a real locally convex space, and $K\subseteq Y$ is a nontrivial cone. Let $\bar{x}\in{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$, i.e., there is $D\in\mathcal{D}(K)$ such that $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,D)$. Define $\bar{D}:=Y\setminus(-\operatorname{int}\,D)$. Assume that $B_{K}$ and $B_{\bar{D}}$ are normlike-bases for $K$ and $\bar{D}$, respectively, constructed with respect to a continuous seminorm $\psi$. Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{\bar{D}}^{0}$ is compact, and
$({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$
Then, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Moreover, any of the equivalent assertions given in Proposition 5.14 is valid.
Proof 5.41.
By Lemma 5.36 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
@@
hence $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,D)\subseteq{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Proposition 5.14 finishes the proof.
Remark 5.42.
Proposition 4.1 gives sufficient conditions for the equality $({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset$, which is given as an assumption in Theorems 5.38 and 5.40. Suppose that there is $(x^{*},\alpha)\in Y^{*}\times\mathbb{P}$ such that (4.4) is valid. Notice that the set $\bar{D}$ is a nontrivial, closed cone in $Y$ with $\bar{D}\cap(-K)=\{0\}$. Assume that the nontrivial, convex cone $K$ is closed and that either ${\mathrm{c}onv}({\mathrm{c}l}(B_{\bar{D}})\cup\{0\})$ is closed or ${\mathrm{c}l}\,B_{\bar{D}}$ is compact. Then, Proposition 4.1 ensures $({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$
Remark 5.43.
The approach in this Section 5.4 follows basically some ideas by Kasimbeyli [45, Th. 4.4 and 5.7], who deals (in a normed setting) with $\varepsilon$-conic neighborhoods for $K$ (as dilating cones) and supposes a separation property for a whole family of dilating cones.
In Section 5.3 one has a separation condition for $A=\mathbb{A}(\bar{x})$ and $-K$ (for a given $\bar{x}$), and one can further ensure that $A$ is convex. Here in Section 5.4 we have a separation condition for $\bar{D}=Y\setminus(-\operatorname{int}\,D)$ and $-K$. One has to notice that the complement cone $\bar{D}$ is usually not convex if $D$ is convex. Moreover, the cone $\mathbb{A}(\bar{x})$, which has a particular structure for specific choices of $\mathbb{A}$, is more easy to handle than the cone $\bar{D}=Y\setminus(-\operatorname{int}\,D)$ based on a general dilating cone $D\in\mathcal{D}(K)$.
@@ 6 Conclusions
It is well-known that several scalarizing functions used in scalarization methods in vector optimization and in the treatment of optimization problems under uncertainty admit certain cone representation and cone monotonicity properties (e.g., methods in the sense of Gerstewitz [20, 21]; Kasimbeyli [45, 46]; Pascoletti and Serafini [53]; Zaffaroni [58]). In the general framework of real topological-linear spaces, our paper provides new insights into the relationships between a vector optimization problem based on an original (convex) cone, an updated vector optimization problem based on an enlarging / dilating cone of Bishop-Phelps type, and scalar optimization problems involving (Bishop-Phelps type) cone representing and cone monotone scalarizing functions. Thus, our new scalarization results with respect to the concepts of weak efficiency and different types of proper efficiency (e.g., in the sense of Benson; Borwein; Henig; Hurwicz) are also useful for several established scalarization methods in vector optimization (see also Remark 5.17).

### v3

@@ 4.5 Scalarization results for the concept of proper efficiency in the sense of Henig
Lemma 35.
Assume that $Y$ is a real locally convex space, $K\subseteq Y$ is a nontrivial cone, and $\bar{D}:=Y\setminus(-\operatorname{int}\,D)$ for some $D\in\mathcal{D}(K)$. Assume that $B_{K}$ and $B_{\bar{D}}$ are normlike-bases for $K$ and $\bar{D}$, respectively, constructed with respect to a continuous seminorm $\psi$. Suppose that ${\rm cl}\,S_{-K}$ is compact, and
$({\rm cl}\,S_{\bar{D}}^{0})\cap({\rm cl}\,S_{-K})=\emptyset.$
@@
Proof.
First, note that $\emptyset\neq K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$ for all $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ taking into account the continuity of $\psi$. By Proposition 26, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
$x^{*}(d)+\alpha\psi(d)\geq 0>x^{*}(-k)+\alpha\psi(k)\quad\mbox{for all }d\in\bar{D}\mbox{ and all }k\in K\setminus\{0\}.$ (25)
Any $x\in C^{>}_{\psi}(x^{*},\alpha)$ satisfies $0>x^{*}(-x)+\alpha\psi(x)$, hence $-x\notin\bar{D}$ in view of (25), and so $x\in\operatorname{int}\,D$. ∎
Theorem 36.
Assume that $Y$ is a real locally convex space, and $K\subseteq Y$ is a nontrivial cone. Let $\bar{x}\in{\rm PEff}_{He}(\Omega\mid f,K)$, i.e., there is $D\in\mathcal{D}(K)$ such that $\bar{x}\in{\rm Eff}(\Omega\mid f,D)$. Suppose that $D$ is closed. Define $\bar{D}:=Y\setminus(-\operatorname{int}\,D)$. Assume that $B_{K}$ and $B_{\bar{D}}$ are normlike-bases for $K$ and $\bar{D}$, respectively, constructed with respect to a continuous seminorm $\psi$. Suppose that ${\rm cl}\,S_{-K}$ is compact, and
$({\rm cl}\,S_{\bar{D}}^{0})\cap({\rm cl}\,S_{-K})=\emptyset.$
Then, the conclusion of Theorem 31 holds.
Proof.
By Lemma 35 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
@@
$\bar{x}\in{\rm Eff}(\Omega\mid f,D)\subseteq{\rm Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha)).$
Finally, noting that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ by Lemma 1 ($7^{\circ}$), Proposition 17 yields the rest of the conclusion. ∎
Theorem 37.
Assume that $Y$ is a real locally convex space, and $K\subseteq Y$ is a nontrivial cone. Let $\bar{x}\in{\rm PEff}_{He}(\Omega\mid f,K)$, i.e., there is $D\in\mathcal{D}(K)$ such that $\bar{x}\in{\rm WEff}(\Omega\mid f,D)$. Define $\bar{D}:=Y\setminus(-\operatorname{int}\,D)$. Assume that $B_{K}$ and $B_{\bar{D}}$ are normlike-bases for $K$ and $\bar{D}$, respectively, constructed with respect to a continuous seminorm $\psi$. Suppose that ${\rm cl}\,S_{-K}$ is compact, and
$({\rm cl}\,S_{\bar{D}}^{0})\cap({\rm cl}\,S_{-K})=\emptyset.$
Then, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that $\bar{x}\in{\rm WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Moreover, if $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing, any of the equivalent assertions given in Proposition 18 is valid.
Proof.
By Lemma 35 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
@@
hence $\bar{x}\in{\rm WEff}(\Omega\mid f,D)\subseteq{\rm WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Proposition 18 finishes the proof. ∎
Remark 23.
Taking Remark 12 into account, in Theorems 36 and 37, we derived necessary (and sufficient) conditions for the proper efficiency in the sense of Henig of $\bar{x}\in\Omega$.
Proposition 19 gives sufficient conditions for the equality $({\rm cl}\,S_{\bar{D}}^{0})\cap({\rm cl}\,S_{-K})=\emptyset$, which is given as an assumption in Theorems 36 and 37. Suppose that there is $(x^{*},\alpha)\in Y^{*}\times\mathbb{P}$ such that (14) is valid. Notice that the set $\bar{D}$ is a nontrivial, closed cone in $Y$ with $\bar{D}\cap(-K)=\{0\}$. Assume that the nontrivial, convex cone $K$ is closed and that either ${\rm conv}({\rm cl}(B_{\bar{D}})\cup\{0\})$ is closed or ${\rm cl}\,B_{\bar{D}}$ is compact. Then, Proposition 19 ensures $({\rm cl}\,S_{\bar{D}}^{0})\cap({\rm cl}\,S_{-K})=\emptyset.$
Remark 24.
The approach in this Section 4.5 follows basically some ideas by Kasimbeyli [8, Th. 4.4 and 5.7], who deals (in a normed setting) with $\varepsilon$-conic neighborhoods for $K$ (as dilating cones) and supposes a separation property for a whole family of dilating cones.
In Section 4.4, one has a separation condition for $A=\mathbb{A}(\bar{x})$ and $-K$ (for a given $\bar{x}$), and one can further ensure that $A$ is convex. Here in Section 4.5 we have a separation condition for $\bar{D}=Y\setminus(-\operatorname{int}\,D)$ and $-K$. One has to notice that the complement cone $\bar{D}$ is usually not convex if $D$ is convex. Moreover, the cone $\mathbb{A}(\bar{x})$, which has a particular structure for specific choices of $\mathbb{A}$, is more easy to handle than the cone $\bar{D}=Y\setminus(-\operatorname{int}\,D)$ based on a general dilating cone $D\in\mathcal{D}(K)$.
@@ 5 Conclusions
It is well-known that several scalarizing functions used in scalarization methods in vector optimization and in the treatment of optimization problems under uncertainty admit certain cone representation and cone monotonicity properties (e.g., methods in the sense of Gerstewitz [26, 27]; Kasimbeyli [8, 9]; Pascoletti and Serafini [10]; Zaffaroni [11]). In the general framework of real topological-linear spaces, our paper provides new insights into the relationships between a vector optimization problem based on an original (convex) cone, an updated vector optimization problem based on an enlarging / dilating cone of Bishop-Phelps type, and scalar optimization problems involving (Bishop-Phelps type) cone representing and cone monotone scalarizing functions. Thus, our new scalarization results with respect to the concepts of weak efficiency and different types of proper efficiency (e.g., in the sense of Benson; Borwein; Henig; Hurwicz) are useful for several established scalarization methods in vector optimization.